# CrystalFlow: A Flow-Based Generative Model for Crystalline Materials

### v1

@@ I Introduction
The prediction of the stable arrangement of atoms within a crystal, given specific chemical compositions and external conditions, is a longstanding challenge known as crystal structure prediction (CSP) [1]. This fundamental problem can be framed as a global optimization task on the potential energy surface (PES) of materials, which has far-reaching implications in the fields of physics, chemistry, and materials science, as the atomic structure of a crystal directly governs its physical and chemical properties [2, 3, 4]. During the past few decades, substantial advancements have been achieved in CSP, primarily attributed to the evolution of a diverse array of CSP methodologies and software, grounded in the integration of sophisticated structural sampling techniques, optimization algorithms, and quantum-mechanical calculations [5, 6, 7, 8, 9, 10]. Today, these computational tools are indispensable in modern computational science, offering critical insights into the phase diagrams of condensed matter and facilitating the design of novel materials with tailored properties. This progress has led to numerous groundbreaking discoveries, such as the identification of high-pressure superhydride superconductors with record-breaking critical temperatures [11, 12, 13, 14].
@@
Recent advancements in deep learning generative models have introduced transformative methodologies for understanding the underlying distributions of high-dimensional data and generating realistic samples. The remarkable capabilities of these generative models have been demonstrated in areas such as large language models [16], image generation[17], and protein design [18]. Coupled with the rapid expansion of comprehensive materials databases, these models present a promising strategy for efficiently sampling the vast crystal space, thereby helping to address the sampling challenges in CSP. Along this line, several generative models specifically designed for crystal structures have been proposed employing variational autoencoders [19, 20, 21, 22, 23, 24], generative adversarial networks [25, 26, 27], diffusion models [28, 29, 30, 31, 32], flow models [33, 34], and autoregressive models [35, 36, 37]. These advancements have led to improved generation performance in terms of metrics such as stability, novelty, and uniqueness of the generated samples, as well as enhanced predictive power.
@@
Most state-of-the-art crystal generative models can be broadly categorized into two main types based on the data representation: graph-based models and string-based models [38]. Graph-based models are frequently combined with diffusion/flow techniques and equivariant message passing networks. This combination is particularly effective in capturing the intrinsic symmetries of crystal structures, such as permutation, rotation and periodic translation. A prominent example in this category is the Crystal Diffusion Variational Autoencoder (CDVAE) [22] which effectively integrates diffusion processes within the framework of a variational autoencoder. By utilizing SE(3) equivariant message-passing neural networks, CDVAE is able to account for key symmetries of crystals, demonstrating superior performance in terms of structural and compositional validity compared to previous models. Recent advancements along this line are represented by models such as DiffCSP [28], MatterGen [29], UniMat [30], and FlowMM [33], which jointly generate lattice, atomic coordinates, and/or atomic types, achieving steady improvements across various standard generative metrics. Furthermore, extensions of CDVAE have been developed for conditional generation based on specific material properties, broadening its applicability [23, 24]. Conversely, string-based models represent crystal structures using standard Crystallographic Information Files (CIFs) or specifically devised string representations, such as SLICES [35]. These models typically employ autoregressive or recurrent neural networks to learn and process the sequential data, capturing the structural information encoded in these formats. Although these models generally do not explicitly consider the symmetry of crystals, they are well-suited for scaling up to work with much larger and more diverse training data, in the same fashion as large language models.
@@
Notably, space group symmetry, a critical inductive bias for modeling crystalline materials, has been introduced into both graph-based and string-based models, such as DiffCSP++ [31], SymmCD [32] and CrystalFormer [36]. This advancement mitigates the challenge of generating structures with high symmetry and enables more data- and compute-efficient learning. Moreover, recent developments have realized the combination of the two types of models, leveraging the strengths of large language models in handling discrete values (atomic types) and flow models in managing continuous values (atomic positions and lattice geometry) [34].
@@ II Method
The CNF can be trained using the CFM [40, 41], a simulation-free method that involves regressing vector fields along fixed conditional probability paths.
Suppose that the marginal probability path $p_{t}(x)$ is a mixture of probatility paths $p_{t}(x|z)$ that vary with some conditioning variable $z$, that is,
@@
generates the probability path $p_{t}(x)$ from initial conditions $p_{0}(x)$.
@@
Here, we utilize the independent coupling variant of the CFM (I-CFM) [43, 44] framework, where the conditioning variable $z$ is defined by the pair $(x_{0},x_{1})$, which represent the initial and terminal points. $p_{t}(x|z)$ is a probability path interpolating between $x_{0}$ and $x_{1}$. The boundary conditions of Eq. 5 satisfy $p_{0}(x|y){\approx}q(x_{0}|y)=q(x_{0})$ being the prior distribution and $p_{1}(x|y){\approx}q(x_{1}|y)$ being target distribution. The prior $q(x_{0})$, conditional probability path $p_{t}(x|z)$ and the conditional vector field $u_{t}(x|z)$ are detailed in Sec. II.3. Then, the training object is
@@
CrystalFlow generates crystalline materials by concurrently evolving the lattice parameters $\mathbf{L}$ and the fractional atomic coordinates $\mathbf{F}$. In the context of crystal generative modeling, it is crucial to incorporate the fundamental periodic-E(3) symmetries of crystalline systems, namely, permutation invariance, rotational equivariance, and periodic translation invariance, to enable data-efficient learning, computational efficiency, and high-quality sampling. To ensure that the push-forward distribution remains invariant under these symmetry transformations, it is generally necessary to employ an equivariant flow in conjunction with an invariant prior with respect to these symmetries [45, 46]. In practice, achieving this invariance can be made more tractable by carefully selecting the representation of the crystal structure $\mathcal{M}$ and designing an appropriate model architecture.
@@
where $\varphi_{k}$, $\varphi_{F}$ and $\varphi_{A}$ are also MLPs.

### v2

@@ I Introduction
The prediction of the stable arrangement of atoms within a crystal, given specific chemical compositions and external conditions, is a longstanding challenge known as crystal structure prediction (CSP) [1]. This fundamental problem can be framed as a global optimization task on the potential energy surface (PES) of materials, which has far-reaching implications in the fields of physics, chemistry, and materials science, as the atomic structure of a crystal directly governs its physical and chemical properties [2, 3, 4]. During the past few decades, substantial advancements have been achieved in CSP, primarily attributed to the evolution of a diverse array of CSP methodologies and software, grounded in the integration of sophisticated structural sampling techniques, optimization algorithms, and quantum-mechanical calculations [5, 6, 7, 8, 9, 10, 11]. Today, these computational tools are indispensable in modern computational science, offering critical insights into the phase diagrams of condensed matter and facilitating the design of novel materials with tailored properties. This progress has led to numerous groundbreaking discoveries, such as the identification of high-pressure superhydride superconductors with record-breaking critical temperatures [12, 13, 14, 15].
@@
Recent advancements in deep learning generative models have introduced transformative methodologies for understanding the underlying distributions of high-dimensional data and generating realistic samples. The remarkable capabilities of these generative models have been demonstrated in areas such as large language models [17], image generation[18], and protein design [19]. Coupled with the rapid expansion of comprehensive materials databases, these models present a promising strategy for efficiently sampling the vast crystal space, thereby helping to address the sampling challenges in CSP. Along this line, several generative models specifically designed for crystal structures have been proposed employing variational autoencoders [20, 21, 22, 23, 24, 25], generative adversarial networks [26, 27, 28], diffusion models [29, 30, 31, 32, 33, 34], flow models [35], and autoregressive models [36, 37, 38, 39, 40, 41, 42]. These advancements have led to improved generation performance in terms of metrics such as stability, novelty, and uniqueness of the generated samples, as well as enhanced predictive power.
@@
Most state-of-the-art crystal generative models can be broadly categorized into two main types based on the data representation: graph-based models and string-based models [43]. Graph-based models are frequently combined with diffusion/flow techniques and equivariant message passing networks. This combination is particularly effective in capturing the intrinsic symmetries of crystal structures, such as permutation, rotation and periodic translation. A prominent example in this category is the Crystal Diffusion Variational Autoencoder (CDVAE) [23] which effectively integrates diffusion processes within the framework of a variational autoencoder. By utilizing SE(3) equivariant message-passing neural networks, CDVAE is able to account for key symmetries of crystals, demonstrating superior performance in terms of structural and compositional validity compared to previous models. Recent advancements along this line are represented by models such as DiffCSP [29], MatterGen [30], UniMat [31], and FlowMM [35], which jointly generate lattice, atomic coordinates, and/or atomic types, achieving steady improvements across various standard generative metrics. Furthermore, extensions of CDVAE have been developed for conditional generation based on specific material properties, broadening its applicability [24, 25]. Conversely, string-based models represent crystal structures using pure textual format [37, 38, 40, 41], such as standard Crystallographic Information Files, or specifically devised ordered-component representations[36, 39, 42], such as SLICES [36]. These models typically employ transformer to learn and process the sequential data, capturing the structural information encoded in these formats. Although these models generally do not explicitly consider the symmetry of crystals, they are well-suited for scaling up to work with much larger and more diverse training data, and enabling co-training with multimodal contents, in the same fashion as large language models.
@@
Notably, space group symmetry, a critical inductive bias for modeling crystalline materials, has been introduced into both graph-based and string-based models, such as DiffCSP++ [32], SymmCD [33], CrystalFormer [39], CHGFlowNet[44], and WyFormer [42]. This advancement mitigates the challenge of generating structures with high symmetry and enables more data- and compute-efficient learning. Moreover, recent developments have realized the combination of the two types of models, leveraging the strengths of large language models in handling discrete values (atomic types) and flow models in managing continuous values (atomic positions and lattice geometry) [38].
@@ II Method
In many scenarios, the vector field $u_{t}(x)$ is intractable, posing significant challenges for analysis and computation. To address this issue, CFM [46, 47] present a simulation-free training strategy that incorporates an additional conditioning variable $z$. This methodolegy is effective provided that a tractable conditional vector field $u_{t}(x|z)$ can be defined. Suppose that the marginal probability path $p_{t}(x)$ is a mixture of probatility paths $p_{t}(x|z)$ that vary with some conditioning variable $z$, that is,
@@
generates the probability path $p_{t}(x)$ from initial conditions $p_{0}(x)$. The CFM training objective becomes $u_{t}(x|z)$.
@@
And the training objective is still $u_{t}(x|z)$.
Here, we utilize the independent coupling variant of the CFM (I-CFM) [49, 50] framework, where the conditioning variable $z$ is defined by the pair $(x_{0},x_{1})$, which represent the initial and terminal points. $p_{t}(x|z)$ is a probability path interpolating between $x_{0}$ and $x_{1}$. The boundary conditions of Eq. 5 satisfy $p_{0}(x|y){\approx}q(x_{0}|y)=q(x_{0})$ being the prior distribution and $p_{1}(x|y){\approx}q(x_{1}|y)$ being target distribution. The prior $q(x_{0})$, conditional probability path $p_{t}(x|z)$ and the conditional vector field $u_{t}(x|z)$ are detailed in Sec. II.3. Finally, the training objective is written as:
@@
CrystalFlow generates crystalline materials by concurrently evolving the lattice parameters $\mathbf{L}$ and the fractional atomic coordinates $\mathbf{F}$. In the context of crystal generative modeling, it is crucial to incorporate the fundamental periodic-E(3) symmetries of crystalline systems, namely, permutation invariance, rotation invariance, and periodic translation invariance, to enable data-efficient learning, computational efficiency, and high-quality sampling. To ensure that the push-forward distribution remains invariant under these symmetry transformations, it is generally necessary to employ an equivariant flow in conjunction with an invariant prior with respect to these symmetries [51, 52]. In practice, achieving this invariance can be made more tractable by carefully selecting the representation of the crystal structure $\mathcal{M}$ and designing an appropriate model architecture.
@@
where $\varphi_{k}$, $\varphi_{F}$ and $\varphi_{A}$ are also MLPs. See Supplementary S6 for other hyperparameters.
@@ Acknowledgements.
Supplementary Information
S1   Symmetry of conditional vector field $u_{t}^{F}$ and conditional probability path $p_{t}^{F}$ for fractional coordinates
Below, we demonstrate that the conditional velocity field $u_{t}^{F}$, as defined in Eq. 12 of the main text, is pairwise invariant under periodic translations. This invariance ensures that the conditional probability path $p_{t}^{F}(\mathbf{F}|z)$, as defined in Eq. 10, is also pairwise invariant under periodic translations.
Specifically, let $\bm{\tau}:=\hat{\bm{\tau}}\mathbf{1}^{\top}$, where $\hat{\bm{\tau}}\in\mathbb{R}^{3\times 1}$ is an arbitrary translation vector and $\mathbf{1}\in\mathbb{R}^{N\times 1}$ is a column vector of ones. The periodic translation operation applied to the fractional coordinates $\mathbf{F}$ is then expressed as $g\circ\mathbf{F}:=w(\mathbf{F}+\bm{\tau})$. We now demonstrate the pairwise periodic translation invariance of $u_{t}^{F}$ under this operation as follows:
$u^{F}_{t}(g\circ\mathbf{F}|g\circ z)$ (S24)
$=$ $w(w(\mathbf{F}_{1}+\bm{\tau})-w(\mathbf{F}_{0}+\bm{\tau})-0.5)-0.5$
$=$ $w(\mathbf{F}_{1}+\bm{\tau}-\lfloor\mathbf{F}_{1}+\bm{\tau}\rfloor-\mathbf{F}_{0}-\bm{\tau}+\lfloor\mathbf{F}_{0}+\bm{\tau}\rfloor-0.5)-0.5$
$=$ $w(\mathbf{F}_{1}-\mathbf{F}_{0}-0.5)-0.5$
$=$ $u^{F}_{t}(\mathbf{F}|z)\;.$
Next, we demonstrate the pairwise periodic translation invariance of the conditional probability path $p_{t}^{F}(\mathbf{F}|z)$, as defined in Eq. 10. Reformulating $p_{t}^{F}(\mathbf{F}|z)$ as:
$p^{F}_{t}(\mathbf{F}|z)=\mathcal{N}_{w}(\mathbf{F}|\mathbf{F}_{0}+tu^{F}_{t}(\mathbf{F}|z),\bm{\sigma}_{\mathbf{F}}^{2}\mathbf{I})\;,$
where $\mathcal{N}_{w}$ denotes the wrapped normal distribution, we proceed with the proof as follows:
$p^{F}_{t}(g\circ\mathbf{F}|g\circ z)$ (S25)
$=$ $\mathcal{N}_{w}(g\circ\mathbf{F};g\circ\mathbf{F}_{0}+tu^{F}_{t}(g\circ\mathbf{F}|g\circ z),\bm{\sigma}_{\mathbf{F}}^{2}\mathbf{I})$
$=$ $\mathcal{N}_{w}(g\circ\mathbf{F};w(\mathbf{F}_{0}+\bm{\tau})+tu^{F}_{t}(\mathbf{F}|z),\bm{\sigma}_{\mathbf{F}}^{2}\mathbf{I})$
$=$ $\mathcal{N}_{w}(w(\mathbf{F}+\bm{\tau});w(\mathbf{F}_{0}+\bm{\tau}+tu^{F}_{t}(\mathbf{F}|z)),\bm{\sigma}_{\mathbf{F}}^{2}\mathbf{I})$
$=$ $\mathcal{N}_{w}(\mathbf{F}+\bm{\tau};\mathbf{F}_{0}+\bm{\tau}+tu^{F}_{t}(\mathbf{F}|z),\bm{\sigma}_{\mathbf{F}}^{2}\mathbf{I})$
$=$ $\mathcal{N}_{w}(\mathbf{F};\mathbf{F}_{0}+tu^{F}_{t}(\mathbf{F}|z),\bm{\sigma}_{\mathbf{F}}^{2}\mathbf{I})$
$=$ $p^{F}_{t}(\mathbf{F}|z)\;.$
Thus, the conditional probability path $p_{t}^{F}(\mathbf{F}|z)$ is shown to be invariant under pairwise periodic translations. This result relies on the properties of the wrapped normal distribution, as established in Lemma 3 of Jiao et al. [29].
S2   Encoding of atom type for de novo generation
In de novo generation tasks, generating the chemical composition $\mathbf{A}$ is a critical step. We employ a one-hot encoding scheme for atomic types to achieve this. To reduce the dimensionality of this representation while maintaining the inherent relationships among chemical species, the periodic table is reorganized into a grid consisting of 13 rows and 15 columns. Each chemical species is assigned a unique position within this grid, defined by specific row and column indices, as illustrated in Table S2–1. These indices are then encoded as one-hot vectors, which are concatenated to form a composite representation for each atomic type: $(\mathbf{a}^{r},\mathbf{a}^{c})$. This approach results in a compact 28-dimensional vector representation for each atomic type, effectively capturing essential chemical relationships in a reduced-dimensional space.
Supplementary Table S2–1: Reorganized periodic table. The elements of sub family, lanthanides and actinides are positioned in the bottom.
0 1 2 3 4 5 6 7 8 9 10 11 12 13 14
0 H He
1 Li Be B C N O F Ne
2 Na Mg Al Si P S Cl Ar
3 K Ca Ga Ge As Se Br Kr
4 Rb Sr In Sn Sb Te I Xe
5 Cs Ba Tl Pb Bi Po At Rn
6 Fr Ra Nh Fl Mc Lv Ts Og
7 Sc Ti V Cr Mn Fe Co Ni Cu Zn
8 Y Zr Nb Mo Tc Ru Rh Pd Ag Cd
9 Hf Ta W Re Os Ir Pt Au Hg
10 Rf Db Sg Bh Hs Mt Ds Rg Cn
11 La Ce Pr Nd Pm Sm Eu Gd Tb Dy Ho Er Tm Yb Lu
12 Ac Th Pa U Np Pu Am Cm Bk Cf Es Fm Md No Lr
S3   Pressure distribution for MP-CALYPSO-60 dataset
Supplementary Fig. S3–1: Distribution of equilibrium pressure in the MP-CALYPSO-60 dataset.
S4   DFT computational details
DFT calculations were performed with the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [59] and all-electron projector-augmented wave method [60], as implemented in the VASP code [56]. An energy cutoff of 520 eV and a Monkhorst-Pack k-point sampling grid spacing of 0.25 Å-1 were used to ensure the convergence of the total energy. The default settings of PBE functional, Hubbard U corrections, and ferromagnetic initializations in Pymatgen [54] MPRelaxSet function were employed, except opting for the W_sv potential for tungsten. The maximum optimization ionic step and the maximum running time are constrained to 150 steps and 20 hours, respectively.
S5   Coverage metric for de novo generation
Following previous works, we employ the coverage metric to measure the structural and compositional similarity between the test set $\mathcal{S}_{t}$ and the generated structures $\mathcal{S}_{g}$. Specifically, let $d_{S}(\mathcal{M}_{1},\mathcal{M}_{2})$ and $d_{C}(\mathcal{M}_{1},\mathcal{M}_{2})$ denote the L2 distances of the CrystalNN [61] structural fingerprints and the normalized Magpie [62] compositional fingerprints. The Coverage Recall is defined as:
[…]
where $\delta_{S}$ and $\delta_{C}$ are pre-defined thresholds. The Coverage Precision is defined similarly by swapping $\mathcal{S}_{t}$ and $\mathcal{S}_{g}$. The recall metrics measure how many ground-truth materials are correctly predicted, while the precision metrics measure how many generated materials are of high quality.
S6   Hyperparameters table for CrystalFlow
Supplementary Table S6–1: Hyperparameters of CrystalFlow model used in this work
   CrystalFlow (CSP)    CrystalFlow (DNG)
Model
Element type encoding dimension $128$ $28$
Gaussian expansion for pressure, number of bases $80$ $80$
Gaussian expansion for pressure, bases start $-2.0$ $-2.0$
Gaussian expansion for pressure, bases stop $5.0$ $5.0$
Time sinusoidal positional encoding dimension $256$ $256$
GNN hidden dimension $512$ $512$
GNN number of layers ($L$) $6$ $6$
Number of frequency for $\mathbf{F}_{ij}^{\text{FT}}$ $256$ $256$
Loss weight $\lambda_{k}$ $10$ $10$
Loss weight $\lambda_{F}$ $1$ $1$
Loss weight $\lambda_{A}$ $10$
Optimizer
Optimizer type Adam Adam
[…]
Learning rate scheduler ReduceLROnPlateau ReduceLROnPlateau
Scheduler patience (epoch) $30$ $30$
Scheduler factor (epoch) $0.6$ $0.6$
[…]
Each element type is encoded by a embedding layer into a vector of length 128 in CSP task, or a concatnated one-hot vector of length 28 in DNG task. Pressures are unit in GPa, and are first standardized on training set, then expanded to 80 equidistant Gaussian bases centered from -2.0 to 5.0 with standard deviation being the interval, i.e., $7/80$ The training batch size is 256 for MP-20 dataset, and 128 for both MPTS-52 and MP-CALYPSO-60 datasets.
S7   Inference time compared to other generative model
Supplementary Table S7–1: The time is evaluated on MP-20 testset and reduced to the generation of 10,000 materials. Best results are marked in bold. Records of FlowMM and FlowLLM are from Ref. [38]
. Device $S$ Time (min/10k) DiffCSP RTX 4090 $1000$ $44.7$ FlowMM A800 $750$ $65.1$ FlowLLM A800 $250$ $89.6$ CrystalFlow RTX 4090 $100$ $4.1$ CrystalFlow RTX 4090 $1000$ $37.0$